# V2I-Calib: A Novel Calibration Approach for Collaborative Vehicle and Infrastructure LiDAR Systems

### v1

@@ Abstract
Cooperative vehicle and infrastructure LiDAR systems hold great potential, yet their implementation faces numerous challenges. Calibration of LiDAR systems across heterogeneous vehicle and infrastructure endpoints is a critical step to ensure the accuracy and consistency of perception system data, necessitating calibration methods that are real-time and stable. To this end, this paper introduces a novel calibration method for cooperative vehicle and road infrastructure LiDAR systems, which exploits spatial association information between detection boxes. The method centers around a novel Overall IoU metric that reflects the correlation of targets between vehicle and infrastructure, enabling real-time monitoring of calibration results. We search for common matching boxes between vehicle and infrastructure nodes by constructing an affinity matrix. Subsequently, these matching boxes undergo extrinsic parameter computation and optimization. Comparative and ablation experiments on the DAIR-V2X dataset confirm the superiority of our method. To better reflect the differences in calibration results, we have categorized the calibration tasks on the DAIR-V2X dataset based on their level of difficulty, enriching the dataset’s utility for future research. Our project is available at https://github.com/MassimoQu/v2i-calib .
@@ I INTRODUCTION
Figure 1: Cooperative Vehicle and Infrastructure LiDAR Calibration Schematic.
To tackle these challenges, the research community has recently seen the emergence of novel approaches. These methods have successfully addressed the issue of time synchronization in collaborative vehicle and infrastructure LiDAR systems [4] [5], ensuring real-time data transmission and the accuracy of data fusion [6]. Despite advancements in spatial alignment, there remain outstanding issues such as the acquisition of spatial initial values [7], the high computational load involved in cross-source point cloud registration [8], or the substantial preliminary costs associated with point cloud mapping techniques [9]. Consequently, achieving spatial alignment in a dynamic environment for vehicle-infrastructure cooperative perception systems while maintaining low computational overhead remains a formidable challenge.
@@
This paper presents a novel calibration method for cooperative vehicle and infrastructure LiDAR systems. The method establishes an affinity matrix for vehicular and roadside nodes based on the Overall IoU metric, enabling the identification of shared matching boxes between vehicle and infrastructure endpoints. Subsequently, it computes and refines the extrinsic parameters associated with matched boxes. The method offers several advantages. Firstly, the proposed method does not require initial extrinsic parameters and meets real-time operation requirements. Secondly, it leverages common target information inherent to traffic scenes, enhancing its generalizability. Furthermore, compared to [10], it exclusively utilizes perception information from target detection, resulting in lower computational complexity and reduced data transmission costs, thereby offering greater potential for practical application. Lastly, V2I-Calib is designed with flexibility, having its components well decoupled, which facilitates easy adaptation to meet specific real-world requirements.
@@
    We introduce the Overall IoU metric, which monitors the real-time calibration performance of extrinsic parameters. Central to this, we propose a method for constructing an affinity matrix for vehicle-road nodes, encoding the correlation between vehicular and roadside targets.
@@
    We propose a calibration method for cooperative vehicle and infrastructure LiDAR systems that thoroughly exploits the spatial associations between detection boxes, characterized by its independence from initial extrinsic parameter values and real-time capabilities.
@@
    The effectiveness of our method is validated through comparative and ablation experiments on the DAIR-V2X dataset, which is also classified according to the difficulty levels of the calibration tasks.
The contents of this paper are organized into the following sections. Initially, in Section II, we will provide a detailed review of the existing vehicle-infrastructure LiDAR calibration methods. Subsequently, in Section III, we will elaborate on the calibration method introduced in this paper, along with its constituent modules. Furthermore, in Section IV, we will conduct experiments to validate the effectiveness of our proposed method. Lastly, the paper will conclude with a summary and future outlook in Section V.
II RELATED WORK
Point cloud registration can be categorized into same-source and cross-source source registration . In essence, the calibration challenge for disparate vehicle and infrastructure LiDAR systems broadly constitutes a cross-source point cloud registration issue [11].
While significant advancements have been made in the registration tasks for same-source point clouds [12] [13] [14] , cross-source point cloud registration in vehicle-infrastructure cooperative systems continues to face challenges. Compared to standard point cloud registration, the task is complicated by a combination of factors such as noise, outliers, density variation, partial overlaps, and scale differences. However, compared to other cross-source point cloud registration tasks[8], vehicle-infrastructure cooperative systems benefit from more consistent and stable information inherent in traffic scenarios, such as elements of traffic participation including pedestrians, vehicles, and traffic signs.
Calibration methods for vehicle-infrastructure LiDAR systems can be categorized based on the source of information used during the calibration process: methods based on high-precision maps and those based on perception results.
High-precision map-based calibration methods align point cloud data collected by vehicle-mounted LiDAR with pre-constructed high-accuracy maps to achieve precise spatial calibration [9] [15]. The advantage of these methods is their utilization of rich environmental information in high-precision maps for stable and reliable calibration. However, they are limited by their reliance on high-quality map data, and the updating and maintenance of these maps pose a challenge. Therefore, these methods are most applicable in relatively stable environments, such as urban arteries and highways, where high-precision maps are available.
Perception result-based calibration methods achieve calibration by identifying and matching prominent features in the environment, such as road markings and traffic signs. These are essentially feature-based calibration methods [16][17], adapted with systematic approaches to feature extraction algorithms known as perception algorithms in the context of autonomous driving. These methods often employ multiple perception algorithms [10] to increase robustness through redundancy but may not perform well in terms of real-time processing. Some approaches combine existing perception algorithms with custom-designed feature extraction modules [18], but the general applicability of these methods across diverse scenarios requires further validation. Other methods attempt to incorporate prior information, such as location, to achieve better spatio-temporal calibration [7], yet their effectiveness at larger data scales remains to be tested.
@@ II RELATED WORK
Figure 2: V2I-Calib Workflow. The method starts with two sets of detection boxes obtained from both vehicle and infrastructure ends. We construct an affinity matrix for vehicle-road nodes using Core affinity and Other affinity. Utilizing matching algorithms, we identify common detection boxes that are matched across the vehicle and infrastructure, after which we proceed to calculate and optimize the extrinsic parameters.
@@ III METHODOLOGY
In this study, we introduce a novel vehicle-infrastructure LiDAR calibration method (V2I-Calib), the overall workflow of which is illustrated in Figure 2. The following sections will first define the problem of extrinsic parameter estimation, and then provide a detailed explanation of the proposed method and its individual components.
III-A Problem Formulation
Traditional point cloud registration methods typically seek a rotation matrix $\mathbf{R}$ and translation vector $\mathbf{t}$ that transform one point cloud $\mathbf{P}_{1}$ onto another $\mathbf{P}_{2}$ by minimizing their discrepancy. Mathematically, this is articulated as finding the minimum of an objective function $E(\mathbf{R},\mathbf{t})$, where $E$ is usually a measure of distance between the point clouds:
$\min_{\mathbf{R},\mathbf{t}}E(\mathbf{P}_{1},\mathbf{R}\mathbf{P}_{2}+\mathbf{t})$ (1)
However, in the context of vehicle-infrastructure (V2I) cooperation, the point clouds from the vehicle and infrastructure inherently exhibit significant discrepancies, to which the traditional methods may not be directly applicable. In response, this paper introduces a detection-box-based method. The main concept is to exploit the spatial information of detection boxes shared between the vehicle and infrastructure.
Specifically, let $\mathbf{B}^{(\text{veh})}$ denote the set of detection boxes from the vehicle, and $\mathbf{B}^{(\text{inf})}$ those from the infrastructure. Our objective is to identify the common detection box set $\mathbf{B}^{(\text{c})}$, comprising the common targets detected by both systems. Mathematically, the common detection box set is expressed as:
$\mathcal{B}^{(c)}=\{\mathbf{B}\mid\mathbf{B}\in\mathcal{B}^{(v)}\cap\mathcal{B}^{(i)}\}$ (2)
After extracting these common detection boxes from the original point clouds, we transform them into abstract point clouds $\mathbf{P}^{(\text{c})}$. This transformation process involves mapping the vertices of detection boxes to point sets, which we denote with the function $\text{Vertices}(\cdot)$. Thus, for each common detection box $\mathbf{B}\in\mathbf{B}^{(\text{c})}$, its corresponding abstract point cloud is defined as:
$\mathbf{P}^{(c)}=\bigcup_{\mathbf{B}\in\mathcal{B}^{(c)}}\text{Vertices}(\mathbf{B})$ (3)
Ultimately, the goal is to obtain the optimal extrinsic parameters $\mathbf{T}$ by minimizing the distance between corresponding abstract point clouds:
$\min_{\mathbf{R},\mathbf{t}}E(\mathbf{P}_{1}^{(c)},\mathbf{R}\mathbf{P}_{2}^{(c)}+\mathbf{t})$ (4)
$T=\begin{bmatrix}\mathbf{R}&\mathbf{t}\\ \mathbf{0}^{T}&1\end{bmatrix}$ (5)
By employing this abstract point cloud registration method based on common detection boxes, we can effectively tackle the challenges of cross-source point cloud registration in V2I cooperative scenarios. The crux of the issue now becomes how to identify the common detection box set $\mathbf{B}^{(\text{c})}$ and then fine-tune the extrinsics, which will be elaborated in subsequent chapters.
III-B Affinity Formulation
In this section, we delve into the process of formulating affinity measures, a pivotal step within the vehicle-infrastructure cooperative LiDAR calibration framework. By crafting a meticulously designed affinity function, we are able to effectively evaluate the congruence between vehicle-side and infrastructure-side detection boxes, providing accurate input data for the subsequent matching module. There are two primary challenges in this process: firstly, the vehicle and infrastructure detection boxes are situated in distinct coordinate systems, complicating the direct utilization of their corresponding geometric relationships; secondly, the randomness inherent in cross-source point cloud perceptions makes distinguishing shared and single-end detection objects in real-world scenarios difficult.
Our work ingeniously addresses these issues through the introduction of a hypothesized matching pairs strategy and the oIoU metric. Below, we will describe the formulation of affinity measures in two parts: core affinity and additional affinities.
III-B1 Core Affinity
@@
The idea behind formulating the Core Affinity is to explore the scene matching scores of the available detection box pairs from both the vehicle and infrastructure. To this end, this section introduces the Overall Intersection over Union (oIoU) as a metric to gauge the degree of scene matching. Here is the mathematical expression:
$\footnotesize\emph{oIoU}=\frac{1}{\max(m,n)}\sum_{i=0}^{\mathbf{m}-1}\sum_{j=0}^{\mathbf{n}-1}\emph{IoU}_{3D}(\mathbf{BBox}_{\emph{infra},i},\mathbf{BBox}_{\emph{vehicle},j})$ (6)
$IoU_{3D(}\mathbf{A},\mathbf{B})=\frac{\text{Vol}(\mathbf{A}\cap\mathbf{B})}{\text{Vol}(\mathbf{A}\cup\mathbf{B})}$ (7)
where $m$ and $n$ represent the number of 3d detection boxes detected by the infrastructure and the vehicle, respectively. $\mathbf{BBox}_{\emph{infra},i}$ and $\mathbf{BBox}_{\emph{vehicle},j}$ are the i-th and j-th 3d detection boxes detected by the infrastucture and the vehicle, respectively. $\emph{IoU}_{3D}(\cdot)$ is the 3d intersection over union function, which calculates the volumetric overlap between 3d detection boxes. $\text{Vol}(\cdot)$ denotes the volume of a 3D detection box, and $\cap$ and $\cup$ represent the intersection and union in the 3D space of the detection boxes, respectively. The complexity of the naive traversal algorithm is $O(m*n)$.
The oIoU metric extends beyond the conventional IoU measure, which traditionally concentrates on the correspondence of individual box pairs. The oIoU fully leverages the positional relationships between 3D detection boxes, innovatively providing a comprehensive assessment of the spatial alignment quality within detection scenarios.
@@
Figure 3: Core Affinity Calculation: two 3D detection boxes are selected from the infrastructure end (red) and the vehicle end (green) as assumed matching pairs to compute the extrinsic parameters at that instance. The entire scene is then transformed and merged to calculate the scene affinity using the overlap IoU (oIoU) metric.
III-B2 Additional Affinity Calculation
@@
III-C Affinity Matrix Formulation and Matching
The primary objective of this module involves constructing an affinity matrix based on the affinity calculation methods, with the aim of identifying the shared matching pairs among vehicle-infrastructure detection boxes.
Our method views the detection boxes as nodes, building weighted graphs $G_{veh}$ and $G_{inf}$ for the vehicle and infrastructure sides, respectively. At this point, the formulation of the affinity matrix and matching issue can be abstracted as a graph matching problem [21]. We can reformulate it into a quadratic assignment problem [22] , described by the vertex affinity matrix $K_{p}$ and the edge affinity matrix $K_{q}$.However, since the quadratic assignment problem is NP-Hard, its computational complexity can easily become a bottleneck in vehicle-infrastructure collaborative scenarios.
We simplify it by integrating the vertex affinity matrix Kp and the edge affinity matrix Kq, reducing it to a linear assignment problem.For each element $K_{p}(i,j)$ , we consider all edges associated with node $i$ and node $j$. Let $E(i)$ be the set of all edges connected to node $i$, and let $K_{q}(u,v)$ denote the affinity between edges $u$ and $v$. The fusion function $F(\cdot)$ integrates the corresponding elements in $K_{q}$ with $K_{p}(i,j)$, for all $u\in E(i)$ and $v\in E(j)$, resulting in the formation of the affinity matrix $A$. The problem is briefly described by the following formula:
$A(i,j)=\emph{F}\left(K_{p}(i,j),\{K_{q}(u,v)|u\in E(i),v\in E(j)\}\right)$ (8)
After obtaining the affinity matrix that reflects the association degree of vehicle-infrastructure nodes, we can employ the Hungarian matching method [23] to find the optimal match within polynomial time. The transition from a quadratic to a linear matching problem might somewhat diminish the ability to independently adjust various affinities. Nonetheless, this simplification aligns well with the needs of our proposed methodology, which merges core and category affinities. Importantly, it significantly lowers the computational complexity of the matching process.
III-D Extrinsics Resolution and Optimization

### v2

@@ Abstract
Cooperative LiDAR systems integrating vehicles and road infrastructure, termed V2I calibration, exhibit substantial potential, yet their deployment encounters numerous challenges. A pivotal aspect of ensuring data accuracy and consistency across such systems involves the calibration of LiDAR units across heterogeneous vehicular and infrastructural endpoints. This necessitates the development of calibration methods that are both real-time and robust, particularly those that can ensure robust performance in urban canyon scenarios without relying on initial positioning values. Accordingly, this paper introduces a novel approach to V2I calibration, leveraging spatial association information among perceived objects. Central to this method is the innovative Overall Intersection over Union (oIoU) metric, which quantifies the correlation between targets identified by vehicle and infrastructure systems, thereby facilitating the real-time monitoring of calibration results. Our approach involves identifying common targets within the perception results of vehicle and infrastructure LiDAR systems through the construction of an affinity matrix. These common targets then form the basis for the calculation and optimization of extrinsic parameters. Comparative and ablation studies conducted using the DAIR-V2X dataset substantiate the superiority of our approach. For further insights and resources, our project repository is accessible at https://github.com/MassimoQu/v2i-calib.
@@ I INTRODUCTION
Figure 1: Cooperative V2I LiDAR Calibration Schematic.
To tackle these challenges, the research community has recently seen the emergence of novel approaches. These methods have provided valuable insights into addressing the issue of time synchronization in collaborative vehicle and infrastructure LiDAR systems [4, 5], contributing to the enhancement of real-time data transmission and the accuracy of data fusion [6]. Despite advancements in spatial alignment, current calibration methods for vehicle-infrastructure cooperative systems typically require initial calibration values as a prerequisite [7, 8], which can be unstable, especially within city canyons. Moreover, existing studies [9] have primarily focused on integrating collaborative calibration methods into cooperative perception strategies, which, to some extent, complicates the model’s interpretability. This approach makes it challenging to quantify whether spatial alignment issues within the calibration process are adequately addressed, which is crucial for further enhancing the overall perception results.
A more focused study on LiDAR calibration in cooperative V2X systems should concentrate on the fundamental issues of point cloud registration. Registration can be classified into same-source and cross-source types based on the similarity of data collection devices, or into local and global types depending on initial calibration values.
In essence, the calibration challenge for disparate vehicle and infrastructure LiDAR systems broadly constitutes a cross-source point cloud registration issue [10]. Furthermore, to maximize the stability of calibration methods within urban canyons, it is advantageous for global point cloud registration to proceed without the necessity for pre-calibration initial values.
Significant progress has been made in same-source point cloud registration [11, 12], but cross-source registration still faces challenges such as noise, outliers, and variations in density and scale [13]. These challenges necessitate complex computational solutions. Yet, vehicle-infrastructure systems benefit from consistent types of traffic participants in their scenarios, providing a stable foundation for data analysis.
Advancements in global point cloud registration, like those proposed by [14, 15, 16], offer initial-calibration-free solutions but demand high computational resources, making real-time processing difficult. Current research is thus directed towards developing efficient algorithms, like graph-based optimization techniques [17], to reduce computational costs.
To sum up, existing approaches must balance the provision of initial extrinsic values with computational efficiency. This trade-off poses significant challenges for the practical deployment of calibration in V2I cooperative perception systems, necessitating innovative solutions that can bridge the gap between theoretical accuracy and real-world applicability.
This paper presents a novel calibration method for V2I cooperative LiDAR systems. The method establishes an affinity matrix for vehicular and roadside nodes based on the innovative Overall Intersection over Union (oIoU) metric, enabling the identification of shared perception targets between vehicle and infrastructure endpoints. Subsequently, it computes and refines the extrinsic parameters associated with matched boxes.
@@
The method offers several advantages. Firstly, the proposed method does not require initial extrinsic parameters and meets real-time operation requirements. Secondly, it leverages common target information inherent to traffic scenes, enhancing its generalizability. Furthermore, compared to [10], it exclusively utilizes perception information from target detection, resulting in lower computational complexity and reduced data transmission costs, thereby offering greater potential for practical application. Lastly, V2I-Calib is designed with flexibility, having its components well decoupled, which facilitates easy adaptation to meet specific real-world requirements.
@@
    We propose a calibration method for cooperative vehicle and infrastructure LiDAR systems that thoroughly exploits the spatial associations between perception objects, characterized by its independence from initial extrinsic parameter values and real-time capabilities.
@@
    We introduce the Overall Intersection over Union metric (oIoU), which monitors the real-time calibration performance of extrinsic parameters. Central to this, we propose a method for constructing an affinity matrix for vehicle-road nodes, encoding the correlation between vehicular and roadside targets.
@@
    The effectiveness of our method is validated through comparative and ablation experiments on the DAIR-V2X dataset.
@@
Figure 2: V2I-Calib Workflow. The process initiates with perception objects from vehicle and infrastructure, forming an affinity matrix using Core and Other affinity. Matching algorithms then identify common objects across both ends, facilitating the calculation and optimization of extrinsic parameters.
II PROBLEM FORMULATION
Traditional point cloud registration methods typically seek a rotation matrix $\hat{\mathbf{R}}$ and translation vector $\hat{\mathbf{t}}$ that transform one point cloud $\mathbf{P}_{1}$ onto another $\mathbf{P}_{2}$ by minimizing their discrepancy. Mathematically, this is articulated as finding the minimum of an objective function $F(\mathbf{P}_{1},\mathbf{P}_{2})$, where $F$ is usually a measure of distance between the point clouds:
$(\hat{\mathbf{R}},\hat{\mathbf{t}})=\arg\min_{\mathbf{R},\mathbf{t}}F(\mathbf{P}_{1}^{(c)},\mathbf{R}\mathbf{P}_{2}^{(c)}+\mathbf{t})$ (1)
However, in the context of vehicle-infrastructure (V2I) cooperation, the point clouds from the vehicle and infrastructure inherently exhibit significant discrepancies, to which the traditional methods may not be directly applicable. In response, this paper introduces an object-aware method. The main concept is to exploit the spatial information of perception objects shared between the vehicle and infrastructure.
Specifically, let $\mathcal{B}^{(\text{veh})}$ denote the set of perception objects from the vehicle, and $\mathcal{B}^{(\text{inf})}$ those from the infrastructure. Our objective is to identify the common perception object set $\mathcal{B}^{(c)}$, comprising the common targets detected by both systems. Mathematically, the common detection box set is expressed as:
$\mathcal{B}^{(c)}=\mathcal{B}^{(veh)}\cap\mathcal{B}^{(inf)}$ (2)
After extracting common perception objects from the original point cloud, we transform these objects into feature point clouds, denoted as $\mathbf{P}^{(\text{c})}$. This transformation involves mapping the vertices of the perception objects into point sets, a process facilitated by the function $\text{Vertices}(\cdot)$. Specifically, the vertices of each detection box $\mathcal{B}_{i}$ are extracted and merged to form the complete feature point cloud:
$\mathbf{P}^{(c)}=\bigcup_{\mathcal{B}_{i}\in\mathcal{B}^{(c)}}\text{Vertices}(\mathcal{B}_{i})$ (3)
This transformation yields a well-characterized feature point cloud, which can then be processed within the framework of classical point cloud registration algorithms, as formulated in Equation (1). Upon obtaining the rotation matrix $\hat{\mathbf{R}}$ and translation vector $\hat{\mathbf{t}}$, these are integrated to produce the optimal extrinsic parameter result $\hat{\mathbf{T}}$.
$\hat{\mathbf{T}}=\begin{bmatrix}\hat{\mathbf{R}}&\hat{\mathbf{t}}\\ \mathbf{0}^{T}&1\end{bmatrix}$ (4)
By employing the point cloud registration method based on common perception objects, we can effectively tackle the challenges of cross-source point cloud registration in V2I cooperative scenarios. The crux of the issue now becomes how to identify the common detection box set $\mathcal{B}^{(c)}$ and then fine-tune the extrinsics, which will be elaborated in subsequent chapters.
@@ III METHODOLOGY
In this study, we introduce a novel vehicle-infrastructure LiDAR calibration method (V2I-Calib), the overall workflow of which is illustrated in Figure 2.
III-A Affinity Formulation
In this section, we explore the development of affinity measures critical to the vehicle-infrastructure cooperative LiDAR calibration framework. By creating a detailed affinity function, we effectively assess the alignment between vehicle-side and infrastructure-side perception objects, essential for the next matching module. The challenges include differing coordinate systems for perception objects and the randomness in cross-source point cloud data, complicating the identification of shared versus unique detection objects. We address these challenges using a hypothesized matching pairs strategy and the Overall Intersection over Union metric(oIoU). We’ll discuss the formulation of affinity measures in two parts: core affinity and additional affinities.
III-A1 Core Affinity
@@
This section outlines the formulation of Core Affinity by evaluating scene matching scores using the Overall Intersection over Union (oIoU) metric, which measures the extent of scene congruence based on detection box pairs from both vehicle and infrastructure sources:
$\emph{oIoU}=\frac{1}{\max(m,n)}\sum_{i=0}^{\mathbf{m}-1}\sum_{j=0}^{\mathbf{n}-1}\emph{IoU}_{3D}(\mathcal{B}^{(inf)}_{i},\mathcal{B}^{(veh)}_{j})$ (5)
$IoU_{3D}(\mathcal{A},\mathcal{B})=\frac{\text{Vol}(\mathcal{A}\cap\mathcal{B})}{\text{Vol}(\mathcal{A}\cup\mathcal{B})}$ (6)
In this model, $m$ and $n$ are the counts of 3D detection boxes by the infrastructure and vehicle, respectively. $\emph{IoU}_{3D}(\cdot)$ calculates volumetric overlaps in 3D space, where $\text{Vol}(\cdot)$ denotes the volumetric overlap between two 3d detection boxes. oIoU advances beyond traditional IoU by incorporating positional relationships among boxes, thus providing a comprehensive evaluation of spatial alignment.
@@
Figure 3: SVD-based extrinsic parameters are computed using selected 3D detection boxes from vehicle (green) and infrastructure (red) as assumed matching pairs [18]. The scene is transformed and merged to assess scene affinity via the oIoU metric.
III-A2 Additional Affinity Calculation
@@
III-B Affinity Matrix Formulation and Matching
This module constructs an affinity matrix from vehicle and infrastructure detection boxes, treating them as nodes within weighted graphs $G_{\text{veh}}$ and $G_{\text{inf}}$. This setup transforms the matching problem into a graph matching problem [20], often modeled as a quadratic assignment problem (QAP) [21], involving vertex affinity matrix $\mathbf{K}_{p}$ and the edge affinity matrix $\mathbf{K}_{q}$.
Due to the NP-Hard nature of QAP and its high computational demand, we simplify the problem to a linear assignment format by integrating $\mathbf{K}_{p}$ and $\mathbf{K}_{q}$ into a single affinity matrix $\mathbf{A}$:
$\mathbf{A}(i,j)=\alpha\cdot\mathbf{K}_{p}(i,j)+\beta\cdot\sum_{u\in E(i),v\in E(j)}\mathbf{K}_{q}(u,v)$ (7)
Here, $E(i)$ represents the set of all edges connected to node $i$. The parameters $\alpha$ and $\beta$ are the adjustment factors for the vertex affinity $\mathbf{K}_{p}(i,j)$ and edge affinity $\mathbf{K}_{q}(i,j)$, respectively, and are empirically set to 0.5. The summation aggregates the edge affinities to complement the vertex affinity in the matrix $\mathbf{A}$.
With affinity matrix $\mathbf{A}$, the Hungarian method [22] efficiently finds optimal matches in polynomial time while reducing computational complexity, suitable for the dynamics of vehicle-infrastructure collaboration.
III-C Extrinsics Resolution and Optimization